OpenStudy (king):

:Help...........for how many positive integers n is n^3 -8n^2+20n-13 a prime number??

OpenStudy (anonymous):

factorize it

OpenStudy (king):

how watever i get is not a prime number

OpenStudy (king):

why?

OpenStudy (king):

oh

OpenStudy (king):

but then ii have to plug in millions of values

OpenStudy (anonymous):

what u want to calculate

OpenStudy (anonymous):

(n^3 -8n^2+20n-13) a prime number

OpenStudy (anonymous):

or n is a prime number

OpenStudy (king):

for n i will have to put millions of values to find how may n positive integers satisfy n^3 -8n^2+20n-13 to be a prime number

OpenStudy (anonymous):

ok

OpenStudy (king):

so,what to do??

OpenStudy (king):

anything i can do to narrow down the possobilities

OpenStudy (king):

no answer given....

OpenStudy (anonymous):

ok see for 2 it is prime number

OpenStudy (anonymous):

wait i took wrong concept

OpenStudy (anonymous):

\[(x-1) \times [(x-7)x + 13]\]

OpenStudy (anonymous):

where x is n

OpenStudy (king):

ok

OpenStudy (anonymous):

now see the first term is (x-1). and other than 2 no other even number is a prime number right

OpenStudy (anonymous):

i got the answer

OpenStudy (king):

no n does not have to be a prime number n^3 -8n^2+20n-13 has to be a prime number

OpenStudy (anonymous):

yeah i know (x−1)×[(x−7)x+13] it has to be prime

OpenStudy (king):

yup so x can be anything greater or equal to 2

OpenStudy (anonymous):

first term is (x-1) and two is the only even prime number.

OpenStudy (anonymous):

but if u will subtract 1 from any other prime number the resultant will be divisible by 2

OpenStudy (anonymous):

getting my point or not

OpenStudy (king):

x can be 4

OpenStudy (anonymous):

yeah u r right

OpenStudy (anonymous):

it means that for any odd integer other than 3 u will not get a prime number

OpenStudy (king):

yeah

OpenStudy (anonymous):

Superb sheggy!

OpenStudy (anonymous):

also the number cannot be graeter than 7

OpenStudy (king):

oh yeah

OpenStudy (anonymous):

thanks fool bhai

OpenStudy (anonymous):

so u are having a limited space to work out that is from 1 to 7

OpenStudy (king):

yes sheg u got it so answer is n=2,3,4

OpenStudy (king):

no actually 2 to 7

OpenStudy (anonymous):

hahahaa right 2 to 7

OpenStudy (anonymous):

but for 7 it will be 0

OpenStudy (king):

wait why cant it be greater than 7?

OpenStudy (anonymous):

try any number u will get the answer

OpenStudy (anonymous):

odd numbers cannot be included ok thats true

OpenStudy (anonymous):

as i had told u any odd number subtracted by one will result in even number so it will be divisible by 2.........do u accept it or not

OpenStudy (king):

is 323 a prime number?

OpenStudy (king):

yeah i acept

OpenStudy (king):

its not div by 17

OpenStudy (anonymous):

323 is divisible by 17

OpenStudy (anonymous):

so 323 is not a prime number

OpenStudy (king):

4387 ?

OpenStudy (anonymous):

what u have put x = ??

OpenStudy (king):

n=20

OpenStudy (king):

1443?

OpenStudy (king):

gtg shall come bak in 15 min

OpenStudy (anonymous):

for n = 20 it should be equal to 5187

OpenStudy (king):

oh yeah 1443?

OpenStudy (king):

no

OpenStudy (anonymous):

20^3 - 8*20^2 + 20*20 -13 =8000 - 3200 + 400 -13 = 8400 - 3213 = 5187

OpenStudy (king):

gtg cum bak in 15 mins

OpenStudy (anonymous):

dear i had told u for no value greater than 7 it would fit

OpenStudy (anonymous):

and both 1443 and 5187 is divisible by 3

OpenStudy (king):

i am bak so what to do now?

OpenStudy (anonymous):

for no value greater than 7 that too even number it will not be a prime number

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=x^3-8x^2%2B20x-13+where+x+%3D+20 visit this website and plug in different value of x u will get the answer

OpenStudy (anonymous):

king u there